OpenStudy (anonymous):

Find derivative of f(x). F(x) = (x^3+x^2-x)/(sqrt(x) +3).

OpenStudy (anonymous):

use quotient rule

OpenStudy (ajprincess):

treat this as u/v form and differentiate.

OpenStudy (anonymous):

Remember that √x = x^½ and 1/√x = x^-½ Also remember that d(x^k)/dx = kx^(k-1) So d(u^½) = ½u^-½ * du/dx (using the chain rule). In this case, u = 5 - 3x and du/dx = -3 d[√(5 - 3x)] = ½/√(5 - 3x) * -3 = -3/[2√(5 - 3x)] ======== Let u = 2 - 9x so du/dx = -9 d[-3(2 - 9x)^(1/3) = d[-3u^(1/3)]/dx d[-3(2 - 9x)^(1/3) = -3d[u^(1/3)]/dx d[-3(2 - 9x)^(1/3) = -3*(1/3)u^(-2/3) * du/dx d[-3(2 - 9x)^(1/3) = -(2 - 9x)^(-2/3) * -9 d[-3(2 - 9x)^(1/3) = 9/(2 - 9x)^(2/3)

hartnn (hartnn):

lol! copy pasted from here... http://answers.yahoo.com/question/index?qid=20081018075010AA7klAr an that too wrong sum... :P

OpenStudy (ajprincess):

Here u=(x^3+x^2-x) and v=(sqrt(x) +3)  Can u do nw?